A rocket of mass 4×10⁴ kg, when full with fuel at launching pad, it's fired vertically upward and a burnout mass of 2×10⁴kg gases are exhausted at the rate of 20 slugs/second, with a velocity of 6×10³ m/s relative to the rocket (exhaust velocity) both quantities are assumed to be constant while the fuel is burning, what's the thrust, neglecting all external forces, what would be the speed of the rocket at burnout?

1
Expert's answer
2022-02-11T15:30:19-0500

"v=v_0\ln (\frac mM)=6.6~\frac{km}s,"

"F=nv_0\frac{\Delta m}{\Delta t}=2.4~GN."